# Ab initio Exact Calculation of Strongly-Correlated Nucleonic Matter

### v2

@@ Abstract
Dense nucleonic matter is of vital importance for understanding compact stars and inferring the transition into deconfined quark phase. We present the ab initio exact calculations of infinite nucleonic matter with the state-of-the-art full configuration-interaction quantum Monte Carlo (FCIQMC) method, enabling us to rigorously benchmark many-body methods and assess the degree to which the nucleonic matter is correlated. This method has been numerically validated by exact diagonalization within a small model space. Calculations of nucleonic matter using chiral nuclear forces reveal that the symmetric nuclear matter is strikingly strongly correlated, raising questions on previous ab initio calculations of nuclear matter with many-body expansion truncations and offering insights into simultaneous descriptions of finite nuclei and infinite nucleonic matter from first principles.
@@
The accurate predictions of equation of state (EoS) of cold dense nucleonic matter that are not accessible by terrestrial experiments are vital for understanding compact neutron stars in the multi-messenger astrophysical era Özel and Freire (2016); Reed et al. (2021); Essick et al. (2021); Neill et al. (2023); Annala et al. (2023). It also provides a stepping stone for inferring the phase transition scenario from nuclear matter to quark matter, informing us of the possible existence of hybrid stars and non-perturbative properties of strong interactions Weih et al. (2020); Zhang and Li (2023); Fujimoto et al. (2023); Huang and others (2022); Kedia et al. (2022). The theoretical challenge is related to the knowledge that nuclear matter is a strongly correlated system, in particular the symmetric nuclear matter (SNM), while the pure neutron matter (PNM) is relatively weakly correlated Gandolfi et al. (2015); Piarulli et al. (2020); Lovato et al. (2022); Drischler et al. (2020). This is indicated by previous studies that different ab initio methods have obvious discrepancies in SNM using hard nuclear forces Carbone et al. (2013); Rrapaj et al. (2016); Hagen and others (2014); Baardsen et al. (2013); Marino et al. (2024); Zhen et al. (2025); Udiani (2025). Therefore, it will be interesting to see how the truncations applied in ab initio calculations of finite nuclei play in infinite nuclear matter.
@@
It is impossible to perform exact diagonalization calculations of infinite nuclear matter by brute force Hjorth-Jensen et al. (2017a). There are ab initio methods with many-body expansion truncations applied to nuclear matter, including the many-body perturbation theory (MBPT) up to the 4th order Tichai et al. (2020); Drischler et al. (2019); Keller et al. (2023), coupled-cluster (CC) truncated at the doubles excitation level with perturbative triple corrections CCD(T) Hagen and others (2014); Baardsen et al. (2013), self-consistent Green’s function (SCGF) theory Rios (2020) based on the algebraic diagrammatic construction (ADC) approximation scheme Raimondi and Barbieri (2018); Barbieri et al. (2022); Marino et al. (2024, 2026), and the in-medium similarity renormalization group (IMSRG) truncated to the normal-ordered two-body level Udiani (2025); Zhen et al. (2025). While these basis-expansion techniques are highly effective with soft potentials, Quantum Monte Carlo (QMC) approaches are generally advantageous for handling harder interactions. In coordinate space, the auxiliary field diffusion Monte Carlo (AFDMC) Wlazłowski et al. (2014); Lynn et al. (2016); Piarulli et al. (2020) and recent variational calculations using neural-network quantum states Carleo and Troyer (2017); Pfau et al. (2020); Hermann et al. (2020); Adams et al. (2021); Keeble and Rios (2020); Luo and Clark (2019) have been developed. In configuration space, methods such as the configuration-interaction Monte Carlo (CIMC) Roggero et al. (2014); Arthuis et al. (2023) are also available. Although these QMC methods can in principle include high-order correlations, they often suffer from the fermion sign problem or rely on the quality of variational ansatz. To address the aforementioned difficulties, we hereby introduce the state-of-the-art full configuration-interaction quantum Monte Carlo (FCIQMC) method for ab initio exact calculations of infinite nuclear matter.
@@
FCIQMC was originally developed in quantum chemistry and quickly has been widely used in molecular and condensed matter physics Booth et al. (2009); Shepherd et al. (2012); Booth et al. (2013); Cleland et al. (2010). FCIQMC has emerged as one of the most accurate many-body methods, particularly for strongly correlated systems Schwarz et al. (2015); Williams and others (2020). Different from post-Hartree-Fock methods, FCIQMC is a projector QMC method that stochastically samples the ground-state wavefunction in the full Hilbert space. This approach also offers distinct advantages over other QMC techniques Spencer and others (2019); Guther and others (2020). Being formulated in a discrete configuration space, unlike coordinate-space methods such as AFDMC, it is not restricted to local potentials and can naturally handle nonlocal forces from chiral effective field theory (EFT). Furthermore, FCIQMC employs a delicate walker annihilation algorithm to overcome the fermion sign problem in QMC methods Booth et al. (2009); Booth and Alavi (2010). This contrasts with methods like AFDMC or CIMC, which require the fixed-node approximation or guiding wavefunctions, introducing a bias that is difficult to assess or eliminate Zhang et al. (1997); Malone et al. (2020); Spencer and others (2019); Guther and others (2020). Therefore, as a method that in-principle is exact and possessing well-controlled systematic errors, FCIQMC holds the promise of providing benchmarking and advancing our understandings of dense nuclear matter and nuclear forces.
@@
where are the Hamiltonian matrix elements and is an energy offset known as the “shift”, which is introduced to prevent the true ground-state from decaying to zero. This master equation is simulated via a stochastic algorithm applied at each time step , which comprises three steps Booth and Chan (2012):
@@
To enhance the efficiency and stability for large systems, we employ the initiator approximation Cleland et al. (2010). This widely-used adaptation designates determinants with a walker population above a certain threshold as “initiators”. Spawning from non-initiator determinants is then restricted to only those determinants that are already occupied. This curtails the proliferation of noise from sparsely populated regions of the Hilbert space. While the initiator approximation introduces a systematic bias, such a bias is progressively reduced by increasing the total number of walkers, allowing for the extrapolation to exact solutions, the unbiased FCI limit Ghanem et al. (2021); Brand et al. (2022). The adaptive-shift method has been used to accelerate the convergence within large model spaces. The full details of the FCIQMC algorithms are shown in the Supplemental Material 1.
@@
Figure 1 presents the correlation energies using different many-body methods and their deviations from the exact solution by direct diagonalzing. As shown, perturbative approaches like MBPT depart from exact solutions rapidly as the interaction strength increases. While non-perturbative methods such as CCD, ADC-SCGF at third-order in the ADC hierarchy plus CCD amplitude corrections [ADC(3)-D] Barbieri and Carbone (2017); Marino et al. (2026), and IMSRG(2) offer significant improvements in the weak-coupling regime, they still exhibit substantial deviations in the strong-coupling regime . In contrast, the FCIQMC results are in remarkable agreement with the exact energies across the entire range of , with statistical errors being almost negligible. The benchmarking shows the superior accuracy of FCIQMC, especially in strongly coupled situations where truncated many-body methods break down (see Fig. 1). Our implementation for the Richardson model is made publicly available for reproducibility and further development Hu .
@@
We now turn to much more challenging calculations of the realistic system of infinite nuclear matter. The many-body Hamiltonian, , includes the kinetic energy term alongside interaction terms corresponding to nucleon-nucleon (NN) and three-nucleon forces (3NF) that are derived from EFT. To compare with other methods, the 3NF is included at the normal-ordered two-body level Hebeler and Schwenk (2010); Hagen and others (2014), although FCIQMC can naturally deal with the full 3NF. The many-body basis is constructed by Slater determinants built with single-particle (SP) momentum eigenstates, on the lattice in momentum space. Our calculations are performed in a cubic box of volume with nucleons at a density , employing periodic boundary conditions which discretize the momenta with a spacing of . The SP basis is defined by a momentum cutoff , which should be sufficiently large to ensure convergence of solutions. Two distinct families of EFT forces at the next-to-next-to-leading-order (N2LO) are used: N2LO(Hüther) Hüther et al. (2020) and Ekström et al. (2018); Jiang et al. (2020), where the later explicitly includes the isobar degree of freedom.
@@
Before proceeding to large-scale calculations with FCIQMC, we perform a rigorous benchmark for symmetric nuclear matter (SNM) in a small model space where exact results can be obtained via direct diagonalization, compared with MBPT and IMSRG(2) methods. The results are shown in Fig. 2. Across all densities and for both families of nuclear forces, FCIQMC calculations with walkers reproduce exact energies with extraordinary accuracy, with deviations being orders of magnitude smaller than other methods. In contrast, both MBPT and IMSRG(2) exhibit significant discrepancies that grow substantially with increasing density and particularly for the harder interactions. MBPT(3) yields reasonable agreement in some cases, which depends on the interaction being used. This benchmark not only validates our FCIQMC implementation for infinite nuclear matter but also provides a clear illustration of the limitations of these widely-used truncated schemes for infinite nucleonic matter.
@@
Next, we present large-scale calculations with the -full interaction, (450 MeV cutoff) Jiang et al. (2020), which has been successful in describing binding energies and radii of finite nuclei up to 132Sn Koszorús and others (2021); Sommer and others (2022); Novario et al. (2023); Nies and others (2023); König and others (2024); Door and others (2025) but whose performance for nuclear matter saturation has been debated Marino et al. (2026); Ekström and others (2024). We compare FCIQMC results with those from other five methods, namely MBPT(2), MBPT(3), CCD, ADC(3)-D and IMSRG(2). All the calculations were performed in the same model space which contains 66 (76) nucleons and 682 (1364) SP states for PNM (SNM). The number of SP states is sufficiently large for the basis convergence. The uncertainty analysis can be found in the Supplemental Material 1.
@@
Acknowledgments.— The authors would like to thank A. Alavi for helpful discussions on the FCIQMC algorithm, as well as J. G. Li, S. Zhang, Carlo Barbieri and Gianluca Colò for useful conversations. This work is supported by the National Key R&D Program of China under Grant No. 2023YFA1606401 and 2024YFA1610900; the National Natural Science Foundation of China under Grants No. 12335007, 12035001, 12475118 and 11921006. We acknowledge the High-Performance Computing Platform of Peking University for providing computational resources. F.M was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 279384907 – SFB 1245, and through the Cluster of Excellence “Precision Physics, Fundamental Interactions, and Structure of Matter” (PRISMA+ EXC 2118/1, Project ID 39083149). We acknowledge the CINECA awards AbINEF (HP10B3BG09) and RespGF (HP10BQMECT) under the ISCRA initiative, for the availability of highperformance computing resources and support. This work used the DiRAC Data Intensive service (DIaL3) at the University of Leicester, managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure.

### v3

@@ Abstract
Dense nucleonic matter is of vital importance for understanding compact stars and inferring the transition into deconfined quark phase. We present ab initio exact calculations of infinite nucleonic matter with the state-of-the-art full configuration-interaction quantum Monte Carlo (FCIQMC) method, enabling us to rigorously benchmark many-body methods and assess the degree to which the nucleonic matter is correlated. This method has been numerically validated by exact diagonalization within a small model space. Calculations of nucleonic matter using chiral nuclear forces reveal that symmetric nuclear matter is strikingly strongly correlated, raising questions on previous ab initio calculations of nuclear matter with many-body expansion truncations and offering insights into simultaneous descriptions of finite nuclei and infinite nucleonic matter from first principles.
@@
The accurate prediction of equation of state (EoS) of cold dense nucleonic matter that is not accessible by terrestrial experiments is vital for understanding compact neutron stars in the multi-messenger astrophysical era Özel and Freire (2016); Reed et al. (2021); Essick et al. (2021); Neill et al. (2023); Annala et al. (2023). It also provides a stepping stone for inferring the phase transition scenario from nuclear matter to quark matter, informing us of the possible existence of hybrid stars and non-perturbative properties of strong interactions Weih et al. (2020); Zhang and Li (2023); Fujimoto et al. (2023); Huang and others (2022); Kedia et al. (2022). The theoretical challenge is related to the knowledge that nuclear matter is a strongly correlated system, in particular the symmetric nuclear matter (SNM), while the pure neutron matter (PNM) is relatively weakly correlated Gandolfi et al. (2015); Piarulli et al. (2020); Lovato et al. (2022); Drischler et al. (2020). This is indicated by previous studies that different ab initio methods have obvious discrepancies in SNM using hard nuclear forces Carbone et al. (2013); Rrapaj et al. (2016); Hagen and others (2014); Baardsen et al. (2013); Marino et al. (2024a); Zhen et al. (2025); Udiani (2025). Therefore, it will be interesting to see how the truncations applied in ab initio calculations of finite nuclei play in infinite nuclear matter.
@@
It is impossible to perform exact diagonalization calculations of infinite nuclear matter by brute force Hjorth-Jensen et al. (2017a). There are ab initio methods with many-body expansion truncations applied to nuclear matter, including the many-body perturbation theory (MBPT) up to the 4th order Tichai et al. (2020); Drischler et al. (2019); Keller et al. (2023), coupled-cluster (CC) truncated at the doubles excitation level with perturbative triple corrections CCD(T) Hagen and others (2014); Baardsen et al. (2013), self-consistent Green’s function (SCGF) theory Rios (2020) based on the algebraic diagrammatic construction (ADC) approximation scheme Raimondi and Barbieri (2018); Barbieri et al. (2022); Marino et al. (2024a, 2026), and the in-medium similarity renormalization group (IMSRG) truncated to the normal-ordered two-body level Udiani (2025); Zhen et al. (2025). While these basis-expansion techniques are highly effective with soft potentials, Quantum Monte Carlo (QMC) approaches are generally advantageous for handling harder interactions. In coordinate space, the auxiliary field diffusion Monte Carlo (AFDMC) Wlazłowski et al. (2014); Lynn et al. (2016); Piarulli et al. (2020) and recent variational calculations using neural-network quantum states Carleo and Troyer (2017); Pfau et al. (2020); Hermann et al. (2020); Adams et al. (2021); Keeble and Rios (2020); Luo and Clark (2019) have been developed. In configuration space, methods such as the configuration-interaction Monte Carlo (CIMC) Roggero et al. (2014); Arthuis et al. (2023) have also been available. Although these QMC methods can in principle include high-order correlations, they often suffer from the fermion sign problem or rely on the quality of variational ansatz. To address the aforementioned difficulties, we hereby introduce the state-of-the-art full configuration-interaction quantum Monte Carlo (FCIQMC) method for ab initio exact calculations of infinite nuclear matter.
@@
FCIQMC was originally developed in quantum chemistry and quickly has been widely used in molecular and condensed matter physics Booth et al. (2009); Shepherd et al. (2012); Booth et al. (2013); Cleland et al. (2010). FCIQMC has emerged as one of the most accurate many-body methods, particularly for strongly correlated systems Schwarz et al. (2015); Williams and others (2020). Different from post-Hartree-Fock methods, FCIQMC is a projector QMC method that stochastically samples the ground-state wavefunction in the full Hilbert space. This approach also offers distinct advantages over other QMC techniques Spencer and others (2019); Guther and others (2020). Being formulated in a discrete configuration space, unlike coordinate-space methods such as AFDMC, FCIQMC is not restricted to local potentials and can naturally handle nonlocal forces from chiral effective field theory (EFT). Furthermore, FCIQMC employs a delicate walker annihilation algorithm to overcome the fermion sign problem in QMC methods Booth et al. (2009); Booth and Alavi (2010). This contrasts with methods like AFDMC or CIMC, which require the fixed-node approximation or guiding wavefunctions, introducing a bias that is difficult to assess or eliminate Zhang et al. (1997); Malone et al. (2020); Spencer and others (2019); Guther and others (2020). Therefore, as a method that in-principle is exact and possessing well-controlled systematic errors, FCIQMC holds the promise of providing benchmarking and advancing our understandings of dense nuclear matter and nuclear forces.
@@
where are the Hamiltonian matrix elements, and is an energy offset known as shift which is introduced to prevent the true ground state from decaying to zero. This master equation is simulated via a stochastic algorithm applied at each time step , which comprises three steps Booth and Chan (2012):
@@
To enhance the efficiency and stability for large systems, we employ the initiator approximation Cleland et al. (2010). This widely-used adaptation designates determinants with a walker population above a certain threshold as “initiators”. Spawning from non-initiator determinants is then restricted to only those determinants that are already occupied. This curtails the proliferation of noise from sparsely populated regions of the Hilbert space. While the initiator approximation introduces a systematic bias, such a bias is progressively reduced by increasing the total number of walkers, allowing for the extrapolation to exact solutions, the unbiased FCI limit Ghanem et al. (2021); Brand et al. (2022). The adaptive-shift method Ghanem et al. (2019); Guther and others (2020) has been used to accelerate the convergence within large model spaces. The full details of the FCIQMC algorithms are shown in the Supplemental Material 1.
@@
Figure 1 presents the correlation energies using different many-body methods and their deviations from the exact solution by direct diagonalzing. As shown, perturbative approaches like MBPT at order (=2, 3, 4) depart from exact solutions rapidly as the interaction strength increases. While non-perturbative methods such as CCD, ADC-SCGF at third-order in the ADC hierarchy plus CCD amplitude corrections [ADC(3)-D] Barbieri and Carbone (2017); Marino et al. (2026), and IMSRG(2) offer significant improvements in the weak-coupling regime, they still exhibit substantial deviations in the strong-coupling regime . In contrast, the FCIQMC results are in remarkable agreement with the exact energies across the entire range of , with statistical errors being almost negligible. The benchmarking shows the superior accuracy of FCIQMC, especially in strongly coupled situations where truncated many-body methods break down (see Fig. 1). Our implementation for the Richardson model is made publicly available for reproducibility and further development Hu .
@@
We now turn to much more challenging calculations of the realistic system of infinite nuclear matter. The many-body Hamiltonian, , includes the kinetic energy term alongside interaction terms corresponding to nucleon-nucleon (NN) and three-nucleon forces (3NF) that are derived from EFT. To compare with other methods, the 3NF is included at the normal-ordered two-body level Hebeler and Schwenk (2010); Hagen and others (2014), although FCIQMC can naturally deal with the full 3NF. The many-body basis is constructed by Slater determinants built from single-particle (SP) momentum eigenstates, on the lattice in momentum space. Our calculations are performed in a cubic box of volume with nucleons at a density , employing periodic boundary conditions which discretize the momenta with a spacing of . The SP basis is defined by a momentum cutoff , which should be sufficiently large to ensure convergence of solutions. Two distinct families of EFT forces at the next-to-next-to-leading-order (N2LO) are used: N2LO(Hüther) Hüther et al. (2020) and Ekström et al. (2018); Jiang et al. (2020), where the latter explicitly includes the isobar degree of freedom.
@@
Before proceeding to large-scale calculations with FCIQMC, we perform a rigorous benchmark for SNM in a small model space where exact results can be obtained via direct diagonalization, compared with MBPT and IMSRG(2) methods. The results are shown in Fig. 2. Across all densities and for both families of nuclear forces, FCIQMC calculations with walkers reproduce exact energies with extraordinary accuracy, with deviations being orders of magnitude smaller than other methods. In contrast, both MBPT and IMSRG(2) exhibit significant discrepancies that grow substantially with increasing density and particularly for the harder interactions. MBPT(3) yields reasonable agreement in some cases, which depends on the interaction being used. This benchmark not only validates our FCIQMC implementation for infinite nuclear matter but also provides a clear illustration of the limitations of these widely-used truncated schemes for infinite nucleonic matter.
@@
Next, we present large-scale calculations with the -full interaction, (450 MeV cutoff) Jiang et al. (2020), which has been successful in describing binding energies and radii of finite nuclei up to 132Sn Koszorús and others (2021); Sommer and others (2022); Novario et al. (2023); Nies and others (2023); König and others (2024); Door and others (2025) but whose performance for nuclear matter saturation has been debated Marino et al. (2026); Ekström and others (2024); Marino et al. (2024b). We compare FCIQMC results with those from other five methods, namely MBPT(2), MBPT(3), CCD, ADC(3)-D and IMSRG(2). All the calculations were performed in the same model space which contains 66 (76) nucleons and 682 (1364) SP states for PNM (SNM). The number of SP states is sufficiently large for the basis convergence. The uncertainty analysis can be found in the Supplemental Material 1.
@@
Acknowledgments.— We are grateful to A. Alavi for valuable suggestions and discussions on the FCIQMC algorithm, and thank J. G. Li, S. Zhang, Carlo Barbieri and Gianluca Colò for useful discussions. This work has been supported by the National Key R&D Program of China under Grant No. 2023YFA1606401 and 2024YFA1610900; the National Natural Science Foundation of China under Grants No. 12335007, 12035001, 12475118 and 11921006, and the High-Performance Computing Platform of Peking University. F.M was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 279384907 – SFB 1245, and through the Cluster of Excellence “Precision Physics, Fundamental Interactions, and Structure of Matter” (PRISMA+ EXC 2118/1, Project ID 39083149). We acknowledge the CINECA awards AbINEF (HP10B3BG09) and RespGF (HP10BQMECT) under the ISCRA initiative, for the availability of highperformance computing resources and support. This work used the DiRAC Data Intensive service (DIaL3) at the University of Leicester, managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure.
Data availability.—The data that support the findings of this article and source code of FCIQMC for the pairing model are openly available Hu .